# Tilus: A Virtual Machine for Arbitrary Low-Precision GPGPU Computation in LLM Serving

### v2

@@ 10. Conclusion
Appendix A Layout Formalization
In this appendix section, we formalize the concept of layout, define primitive layouts, composition, division operators, and layout broadcast. Furthermore, we prove that the set of layouts with the composition operator forms a monoid.
Definition 0 (Layout).
A layout $f$ is a function that maps a pair of thread index $t$ and local element index $i$ to the logical index $I$ of the element in the tensor:
[…]
where $T$ is the number of threads, $m$ is the number of elements each thread holds, $r$ is the rank of the tensor, $\mathbb{N}_{n}=\{0,1,\dots,n-1\}$ and $S=(n_{1},n_{2},\dots,n_{r})$ is the shape of the tensor.
Definition 0 (Local Layout).
Given an $r$-rank tensor with shape $(n_{1},n_{2},\dots,n_{r})$, we define the local layout as a layout with a single thread ($T=1$), where all elements of the tensor are stored in row-major order:
$f(t,i)=(i/(n_{2}n_{3}\dots n_{r}),\;i/(n_{3}n_{4}\dots n_{r})\%n_{2},\;\dots,\;i\%n_{r}).$
Here, $0\leq t<1=T$ and $0\leq i<n_{1}n_{2}\dots n_{r}=m$. In code, we represent such a layout as local(n1, n2, ..., nr).
Definition 0 (Spatial Layout).
Given an $r$-rank tensor with shape $(n_{1},n_{2},\dots,n_{r})$, we define the spatial layout as a layout that distributes all elements across $n_{1}n_{2}\dots n_{r}$ threads, where each thread holds a single element:
$f(t,i)=(t/(n_{2}n_{3}\dots n_{r}),\;t/(n_{3}n_{4}\dots n_{r})\%n_{2},\;\dots,\;t\%n_{r}).$
Here, $0\leq t<n_{1}n_{2}\dots n_{r}=T$ and $0\leq i<1=m$. In code, we represent such a layout as spatial(n1, n2, ..., nr).
Definition 0 (Layout Composition).
Given two layouts $f$ and $g$ with the same rank $r$, we define the composition layout $h$ of $f$ and $g$ as:
$h(t,i)$ $=f(t\;/\;T_{g},i\;/\;m_{g})\odot S_{g}+g(t\;\%\;T_{g},i\;\%\;m_{g}),$
$T_{h}$ $=T_{f}T_{g},$
$m_{h}$ $=m_{f}m_{g},$
$S_{h}$ $=S_{f}\odot S_{g},$
where $T_{h,f,g}$, $m_{h,f,g}$, and $S_{h,f,g}$ are the number of threads, local elements per thread, and shape of layouts $h,f,g$, respectively. Here, $\odot$ represents elementwise multiplication. The composition is denoted as $h=f\circ g$. In code, we use spatial(2, 3).local(3, 2) to denote the composition of spatial(2, 3) and local(3, 2).
Theorem 5 (Layout Composition Associativity).
The layout composition operation is associative:
$f\circ(g\circ h)=(f\circ g)\circ h.$
Proof.
Let $f,g,h$ be layouts with the following properties:
$f(t,i):\mathbb{N}_{T_{f}}\times\mathbb{N}_{m_{f}}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}},$
$g(t,i):\mathbb{N}_{T_{g}}\times\mathbb{N}_{m_{g}}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}},$
$h(t,i):\mathbb{N}_{T_{h}}\times\mathbb{N}_{m_{h}}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}}.$
For $f\circ g$, the layout $k(t,i)$ is:
$k(t,i)=f(t/T_{g},i/m_{g})\odot S_{g}+g(t\%T_{g},i\%m_{g}),$
where $T_{k}=T_{f}T_{g},\;m_{k}=m_{f}m_{g},\;S_{k}=S_{f}\circ S_{g}$.
For $g\circ h$, the layout $l(t,i)$ is:
$l(t,i)=g(t/T_{h},i/m_{h})\odot S_{h}+h(t\%T_{h},i\%m_{h}),$
where $T_{l}=T_{g}T_{h},\;m_{l}=m_{g}m_{h},\;S_{l}=S_{g}\circ S_{h}$.
Now consider $f\circ(g\circ h)$. we get layout $p$:
$p(t,i)=f(t/T_{l},i/m_{l})\odot S_{l}+l(t\%T_{l},i\%m_{l}).$
Expanding $l(t\%T_{l},i\%m_{l})$:
$l(t\%T_{l},i\%m_{l})=$ $g((t\%T_{l})/T_{h},(i\%m_{l})/m_{h})\odot S_{h}$
$+h((t\%T_{l})\%T_{h},(i\%m_{l})\%m_{h}).$
Substituting this into $p(t,i)$:
$p(t,i)=$ $f(t/T_{l},i/m_{l})\odot S_{l}$
$+g((t\%T_{l})/T_{h},(i\%m_{l})/m_{h})\odot S_{h}$
$+h((t\%T_{l})\%T_{h},(i\%m_{l})\%m_{h}).$
For $(f\circ g)\circ h$, we get layout $q$:
$q(t,i)=k(t/T_{h},i/m_{h})\odot S_{h}+h(t\%T_{h},i\%m_{h}).$
Expanding $k(t/T_{h},i/m_{h})$:
$k(t/T_{h},i/m_{h})=$ $f((t/T_{h})/T_{g},(i/m_{h})/m_{g})\odot S_{g}$
$+g((t/T_{h})\%T_{g},(i/m_{h})\%m_{g}).$
Substituting this into $q(t,i)$:
$q(t,i)=$ $f(t/T_{l},i/m_{l})\odot S_{l}$
$+g((t\%T_{l})/T_{h},(i\%m_{l})/m_{h})\odot S_{h}$
$+h((t\%T_{l})\%T_{h},(i\%m_{l})\%m_{h}).$
Since $p(t,i)=q(t,i)$, we conclude:
$f\circ(g\circ h)=(f\circ g)\circ h.$
This completes the proof. ∎
Definition 0 (Layout Broadcast).
To handle layouts of different ranks, we define the layout broadcast operation, which increases the rank of a layout by prepending zeros to the output.
Let $f(t,i):\mathbb{N}_{T}\times\mathbb{N}_{m}\to\mathbb{N}_{n_{1}}\times\cdots\times\mathbb{N}_{n_{r}}$ be a layout of rank $r$. To broadcast $f$ to rank $r^{\prime}>r$, we define:
$f^{\prime}(t,i)=(0,\dots,0,f(t,i)),$
where $r^{\prime}-r$ zeros are prepended. In code, this is denoted as broadcast(f, rank).
Using broadcast, we can compose any two layouts $f$ and $g$ of ranks $r_{f}$ and $r_{g}$, respectively, by broadcasting the smaller rank to match the larger one before composition:
$f\circ g=\text{broadcast}(f,\max(r_{f},r_{g}))\circ\text{broadcast}(g,\max(r_{f},r_{g})).$
Definition 0 (Layout Division).
Given two layouts $f$ and $g$, if there exists a layout $h$ such that $f=h\circ g$, we define $h$ as the result of dividing $f$ by $g$:
$h=f/g.$
In code, this is denoted as f/g.
Figure 14. The optimized matrix multiplication with proper tile sizes and software pipeline optimization.
Theorem 8 (Monoid Structure of Layouts).
The set of layouts $L$, under the composition operator $\circ$, forms a monoid.
Proof.
To prove that $(L,\circ)$ is a monoid, we verify the three properties:
  • Clousure: The composition of two layouts is a layout, as shown in the definition of composition.
  • Associativity: This was proven earlier in the theorem on layout composition associativity.
  • Identity Element: Define the identity layout $e$ as local(1) (or equivalently spatial(1)). It satisfies $f\circ e=e\circ f=f$ for any layout $f$.
Thus, $(L,\circ)$ satisfies the monoid properties. ∎
Appendix B Optimized Matrix Multiplication
In Figure 2 of the main paper, we presented an example of quantized matrix multiplication (FP16xINT6). This example primarily demonstrates the programming of the virtual machine using algebraic layouts, thread-block-level instructions, and the handling of low-precision data types. However, it does not incorporate optimized tiling sizes or advanced techniques such as software pipelining (Huang et al., 2023).
In this section, we extend the example from the main paper by introducing software pipelining and selecting better tiling sizes to program an efficient matrix multiplication. The complete program is illustrated in Figure 14.
B.1. New Optimizations
The key improvement in this program compared to the simplified version in Figure 2 is the use of shared memory to store global sub-tensors, which significantly enhances data reuse since shared memory is an order of magnitude faster than global memory (NVIDIA Corporation, 2024a). Shared memory is also employed as a staging area for results before writing them back to global memory, ensuring coalesced global memory access (NVIDIA Corporation, 2024a). Asynchronous data transfer from global memory to shared memory is achieved using the CopyAsync, CopyAsyncCommitGroup, and CopyAsyncWaitGroup instructions. To maximize performance, memory loading is pipelined with computation. This is accomplished by allocating three shared memory buffers for each operand (A and B). During the main loop (Line 31-45), the shared memory buffer for the next stage (preload_stage) is populated (Line 39-40) while computations are performed using the buffer for the current stage (current_stage). These optimizations are essential to achieve state-of-the-art performance. Our virtual machine not only supports these optimizations seamlessly but also accommodates arbitrary low-bit-width quantized data types via register tensor reinterpretation (Line 36).
B.2. Line-by-Line Code Analysis
Initialization: Workload sizes, tiling sizes, and the number of pipeline stages are defined (Line 1-3). Register and shared memory tensor layouts for the virtual machine program are then specified (Line 5-11). There are two new operations over the layouts that we did not introduce in the main text: reduce(...) and swizzle(...). The reduce(...) layout reduces a layout over some dimensions. If there are multiple indices over that dimension are stored in the same thread, the thread will only keep one copy of the elements. If there are multiple threads along the reduction dimension, we will retain one copy of the element for each thread. In this example, reduce(spatial(1, 1, 4), dims=[2]) is a layout with shape $(1,1)$ and $4$ threads. Each thread will hold the element $(0,0)$. On the other hand, the swizzle operator swizzle a layout by performing the following operation:
$\text{swizzle}(i,j)=$ $\left(i,j\oplus\left\lfloor\frac{i}{2^{\text{log\_step}}}\right\rfloor\right)$
$g(t,i)=$ $\text{swizzle}(f(t,i))$
where $\oplus$ is bitwise-xor operation, and log_step is the parameter of the swizzle operator. The dim=1 in the code indicates that we hope to swizzle the dimension $1$, that’s why we changed the dimension $1$ and keep the other dimension unchanged. The swizzle operation is necessary to avoid the shared memory bank-conflict (NVIDIA Corporation, 2024a) for both the global to shared memory transfer (via CopyAsync) and the shared memory to register transfer (via LoadShared).
Global Tensor Views and Memory Allocation: Global tensor views and offsets for thread blocks are initialized (Line 14-18). Shared memory for software pipelining and registers for accumulation are allocated (Line 19-21).
Preloading Shared Memory: The first STAGES - 1 iterations preload the shared memory buffers (Line 22-28). Two variables, current_stage and preload_stage, track the buffers used for computation and memory copying (Line 29-30).
Main Loop: For each iteration, the k-dimension is further tiled to reduce register usage (Line 33-38). Shared memory buffers are loaded into register tensors, reinterpreted for quantized data types, and cast to the computation data type (e.g., FP16). The dot product operation is then performed. Simultaneously, asynchronous global-to-shared memory transfers are issued using a separate buffer (Line 39-44). This multi-buffer approach enables parallel utilization of GPU computation and memory transfer units.
Result Storing: After completing the main loop, the accumulation tensor is cast from float32 to float16 (Line 46). The tensor is stored in shared memory (Line 48), reloaded into a new register tensor (Line 50), and written back to global memory with coalesced access (Line 51).
This improved program illustrates the flexibility of the virtual machine in implementing advanced optimizations while supporting low-precision data types efficiently. By utilizing shared memory, asynchronous transfers, and software pipelining, it achieves significant performance gains.
Appendix C Program Compilation and Runtime
We provides a domain-specific language (DSL) for our virtual machine in Python to allow programmers to directly write the virtual machine programs in Python, making it easy to integrate the produced kernels with the rich deep learning ecosystem in Python. Given a program, we take several steps to compile it to GPU executable code.
Step 1: Global and Shared Memory Planning Each GPU kernel could use some launch-time known size of shared memory space. To simplify GPU programming, we allow the users to allocate shared memory multiple times in the program on demand. Thus, we need a shared memory planner to calculate the size of shared memory the virtual machine needs, and map the shared tensor to one region of the kernel’s shared memory space. Similar to shared memory planning, we also require a global memory planner to manage the allocation of global memory shared by all thread blocks. We will request the runtime system of our virtual machine to allocate a workspace in global memory, enabling the kernel to use this workspace during its execution.
Step 3: Code Emitting for Each Instruction. We emit low-level GPU code for each VM instruction one by one. In our implementation, we use the Hidet IR (Ding et al., 2023) to represent the low-level GPU code. During this process, we will do instruction selection to select the most efficient low-level instructions when feasible. For example, we use lds PTX instruction (NVIDIA Corporation, 2024b) to load the data from shared memory to register. However, a more efficient PTX instruction ldmatrix could also be used if the layout of the loaded register tensor is divisible by layout spatial(8, 4).repeat(1, 4). Besides, we also try to perform automatic vectorization for memory loading and storing instructions. For example, we will try to use the vectorized instructions like cp.async.v4, lds128 and ldg128 to maximize the memory accessing efficiency.
Step 4: Lowering Low-Precision Data Types. After we emit thread-block-level instructions to the low-level IR, we will apply the passes that implement the the rules discussed in Section 7 to transform all low-precision operations in the low-level IR to corresponding operations on hardware-friendly types. en most cases, only the vectorized type casting from low-precision type to standard type (e.g., float16) will be applied since the memory loading of low-precision data will be replaced by standard types thanks to our layout formalization and register tensor reinterpretation. Afterwards, we generate the CUDA (for NVIDIA GPUs) code from the low-level IR, and finally use the compiler nvcc to compile the source code to hardware binary.
Vitual Machine Runtime. The compiled binary could be loaded by the virtual machine runtime. The runtime also maintains internal states to serve the kernel execution: 1) an on-demand allocated workspace memory that can be requested by the compiled kernels; 2) an execution context that stores the CUDA stream that the kernel will be launched on; 3) the kernels cached in memory.
Appendix D Profiling of Quantized Kernels
Figure 15. GPU hardware utilization for cuBLAS, Ladder, and Tilus kernels. The figure details the throughput utilization of memory units and pipeline utilization of computation units. Nsight Compute is used to collect the results.
We use Nsight Compute to profile the kernels of cuBLAS, Ladder, and Tilus, with results shown in Figure 15. The figure details GPU hardware utilization, dividing it into throughput utilization for L1 Cache, L2 Cache, and DRAM (first three columns) and computation unit utilization for ALU, FMA, and Tensor Core (last three columns). Specifically, ALU handles bit manipulation and logical instructions, FMA performs fused multiply-add operations for integers and floating-point numbers, and Tensor Core represents the utilization of NVIDIA’s specialized Tensor Core pipeline (NVIDIA Corporation, 2024a). From row B of Figure 15, Ladder kernels primarily rely on CUDA Cores, with $40.5\%$ FMA utilization and no Tensor Core usage. However, the kernel suffers from significant DRAM bandwidth limitations ($96.6\%$ utilization), as each thread independently loads operand A from global memory without exploiting shared memory for data reuse. This inefficiency worsens with larger batch sizes, as seen in Figure 13, where performance degrades for batch sizes of 1, 4, and 8 (Wang et al., 2024). For batch sizes of 16 or more, Ladder transitions to Tensor Core usage (row C, Figure 15). Despite this transition, Ladder does not utilize cp.async for asynchronous memory transfers or software pipelining to overlap computation and memory operations. These limitations hinder efficient resource utilization. Additionally, Ladder’s Tensor Core execution strategy is flawed; padding in the batch size dimension results in tensor core instructions up to eight times more than necessary. In contrast, Tilus overcomes these inefficiencies using an algebraic layout system. It interprets quantized weights with standard data types and employs cp.async to optimize memory access patterns. Moreover, Tilus integrates software pipelining to overlap memory transfers and computation, leading to superior performance, as evident in row D of Figure 15.

### v3

@@ Acknowledgement
Appendix A Artifact Appendix
A.1. Abstract
We provide artifacts to reproduce all experimental results discussed in the evaluation section. These artifacts include the compiler implementation, kernels in our domain-specific languages, and scripts for running experiments and plotting figures. To simplify reproduction, all necessary dependencies are packaged within a Docker image. A Dockerfile is also included to demonstrate the image building process. We do not provide the models directly; instead, they and their metadata will be automatically downloaded from Hugging Face Hub when the experiment scripts are launched.
A.2. Artifact check-list (meta-information)
  • Model: Gemma, Llama, Qwen
  • Run-time environment: Linux, CUDA
  • Hardware: NVIDIA L40s
  • Metrics: Latency (ms), Speedup
  • Output: Efficient kernels, Numerical results
  • Experiments: Automated scripts in docker
  • How much disk space required (approximately)?: 25 GiB
  • How much time is needed to prepare workflow (approximately)?: 10 minutes
  • How much time is needed to complete experiments (approximately)?: 3 hours
  • Publicly available?: Yes
  • Code licenses (if publicly available)?: Apache 2.0
  • Workflow automation framework used?: Docker
  • Archived (provide DOI)?:
    https://doi.org/10.5281/zenodo.16756859
A.3. Description
A.3.1. How to access
We have open-sourced our artifacts at https://github.com/yaoyaoding/tilus-artifacts. To pull the Docker image and perform experiments, follow the guide in the README.md file. The code itself is several megabytes, while the Docker image is approximately 21 GiB. We only fetch model meta-information (e.g., number of layers, layer size) from Hugging Face Hub, and dummy weights are used; therefore, the models do not consume significant disk space.
A.3.2. Hardware dependencies
Our experiments were primarily conducted on NVIDIA L40s. To perform a hardware ablation study, we also ran some experiments on NVIDIA A100 and NVIDIA H100. Any NVIDIA GPU with compute capability $\geq 8.0$ should be able to run our artifacts and observe speedup, though the specific numbers might vary slightly.
A.3.3. Software dependencies
We provide a Docker image with all software dependencies pre-installed. Therefore, only the following software is required:
  • NVIDIA GPU driver $\geq$ 565.57.01
  • NVIDIA container toolkit
  • Docker
We have the following packages pre-installed:
  • PyTorch v2.5.1
  • Triton v3.1.0
  • BitBLAS v0.0.1.dev15
  • Marlin v0.1.1
  • vLLM 0.7.3
A.3.4. Data sets
We use dummy inputs and weights because we are solely focused on system performance, which is independent of the input and weight content.
A.3.5. Models
The artifact uses three models for end-to-end evaluation: Gemma-2-9B, QWen-2.5-32B, and Llama-3.3-70B. The meta-information of these models will be automatically fetched from Hugging Face Hub (some may require a Hugging Face token).
A.4. Installation
First, clone the artifact Git repository:
git clone https://github.com/yaoyaoding/tilus-artifacts.git tilus
Then, install Docker and the NVIDIA Container Toolkit by following the README.md in the artifact.
A.5. Experiment workflow
All experiments can be executed with:
bash run.sh
This command will create a Docker container and sequentially run all experiments within it.
A.6. Evaluation and expected results
Upon completion of the experiments, a folder named results or precompiled-results will be created under the artifact directory. This folder will contain four figures of the evaluation results, corresponding to those presented in the evaluation section.
A.7. Methodology
Submission, reviewing and badging methodology: